OpenStudy (anonymous):

Fifteen more than four times a number is 6 more than five times the number. Find the number.

OpenStudy (anonymous):

21/9 -9 -9/21 2 9

OpenStudy (anonymous):

15 + 4x = 6 + 5x x = 9

OpenStudy (anonymous):

ur answer is e, also let me know if you understand how i got my equation